OpenStudy (babynini):

Particles moving on vertical lines! Please check my work and help me with what to do next :)

OpenStudy (babynini):

Picture 1: original question picture 2: my work work so far o.o

OpenStudy (babynini):

@ganeshie8 :)

OpenStudy (babynini):

@zepdrix ^-^

OpenStudy (anonymous):

a is correct. Well done! (still looking through the others)

OpenStudy (babynini):

Thank you! I'll brb.

OpenStudy (anonymous):

In part b, when you factor the polynomial, be careful of the signs in the terms. Double check that part.

OpenStudy (babynini):

hmm

OpenStudy (babynini):

wait how can I factor that o.o

OpenStudy (babynini):

without making the 6 positive

jimthompson5910 (jim_thompson5910):

You'll have to use the quadratic formula. It doesn't factor.

jimthompson5910 (jim_thompson5910):

For part (b), you will follow these steps step 1) solve v(t) = 0 by using the quadratic formula step 2) make a sign chart to figure out the intervals where v(t) > 0 (when the particle is moving upward) and where v(t) < 0 (when the particle is moving downward). Keep in mind that t >= 0

OpenStudy (babynini):

Sorry, you guys! I had a phone call o.o but I am back now :)

OpenStudy (babynini):

Gah quadratic formula is too long ago -.- how do I do that?

jimthompson5910 (jim_thompson5910):

Quadratic Formula \[\Large x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

jimthompson5910 (jim_thompson5910):

In the case of v(t) = 2t^2 - 14t - 12, we have a = 2 b = -14 c = -12

OpenStudy (babynini):

ok, give me a moment :)

OpenStudy (babynini):

(7- sqroot73)/2 (7+sqroot73)/2

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

what are those values approximately equal to? in decimal form

OpenStudy (babynini):

- one = -0.77200 + one = 7.77200

jimthompson5910 (jim_thompson5910):

so we can ignore the root `t = -0.77200` since `t >= 0`

jimthompson5910 (jim_thompson5910):

we only have to deal with `t = 7.77200` now make a sign chart with 7.77200 on the number line and determine where v(t) is positive or negative

OpenStudy (babynini):

by choosing numbers on either side of it, yes?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (babynini):

It's positive when it's larger than 77.772

jimthompson5910 (jim_thompson5910):

I think you mean 7.772, but yes, correct

jimthompson5910 (jim_thompson5910):

the object is moving upward when t > 7.772

OpenStudy (babynini):

and downward when t<7.772

jimthompson5910 (jim_thompson5910):

when 0 <= t < 7.772

OpenStudy (babynini):

is that equals meant to be in there?

jimthompson5910 (jim_thompson5910):

yes because \(\Large t \ge 0\)

jimthompson5910 (jim_thompson5910):

so \(\Large 0 \le t < 7.772\)

OpenStudy (babynini):

Oh I see :) thanks. So I guess that answers of all b!

OpenStudy (babynini):

To answer c?

OpenStudy (babynini):

I plug 0 into the original function y(x) to find where the particle begins?

jimthompson5910 (jim_thompson5910):

let f(x) be the position function on a vertical number line f(0) is the object's position at time t = 0 seconds f(5) is the object's position at time t = 5 seconds |f(5) - f(0)| is the total distance the object travels from t = 0 to t = 5 seconds keep in mind that if the object turned around anywhere in the interval [0,5], then we'd have to break up the interval into smaller pieces. But the object doesn't turn until 7.772 seconds, which is outside the interval [0,5]

OpenStudy (babynini):

So in this case I can just plug 0 and 5 into the original function.

jimthompson5910 (jim_thompson5910):

btw that's equivalent to saying \[\Large \int_0^5 v(t)dt\]

jimthompson5910 (jim_thompson5910):

yes and subtract. Make the final result positive

OpenStudy (babynini):

Emm.. y(0)=12 y(5)=-139.667

jimthompson5910 (jim_thompson5910):

both incorrect

OpenStudy (babynini):

Ai. Do I plug it into v(t)??

jimthompson5910 (jim_thompson5910):

into the position function, not the velocity function

OpenStudy (babynini):

Yeah that's what I did and got the first values? (2/3)(5)^3-7(5)^2-12(5)+12

jimthompson5910 (jim_thompson5910):

I see `+2` at the end, not `+12`

OpenStudy (babynini):

oh gosh. I copied it down wrong on my paper.

OpenStudy (babynini):

after subtracting we get 151.6667

jimthompson5910 (jim_thompson5910):

me too

OpenStudy (babynini):

Excellent. Sorry about the confusion. Part d!

OpenStudy (babynini):

So it speeds up when acceleration and velocity match signs (both - or both +) and slows down when they are not matching. Yes?

jimthompson5910 (jim_thompson5910):

Q: When is the particle speeding up? When is it slowing down? A: If the velocity and acceleration are the same sign (both positive together or both negative together), then the object is speeding up. Otherwise, if the velocity and acceleration differ in sign, then the object is slowing down. See the attached image chart

jimthompson5910 (jim_thompson5910):

Yes you beat me to it

OpenStudy (babynini):

haha. So where do I test if acceleration is equal (Signs) to velocity?

OpenStudy (babynini):

the same as I did for finding when the particle was moving upward or downward?

jimthompson5910 (jim_thompson5910):

I would make a sign chart for both v(t) and a(t). Line up the number lines (one over the other)

OpenStudy (babynini):

 This is pretty much what i did for velocity. I can plug those random values into a(t) as well and see where the signs match?

jimthompson5910 (jim_thompson5910):

this is how I did it (using geogebra and MS paint)

jimthompson5910 (jim_thompson5910):

I found that a(t) = 0 leads to 4t - 14 = 0 which means t = 14/4 = 7/2 = 3.5 that's when a(t) flips from negative to positive

OpenStudy (babynini):

So it's speeding up at: (0,3.5)u(7.5,10) and slowing down at (3.5,7.5)

jimthompson5910 (jim_thompson5910):

where are you getting 7.5 ?

OpenStudy (babynini):

*7.772 I was just rounding. sorry.

OpenStudy (babynini):

since 10 is not a set number, should I put to infinity? or is that not safe to say?

jimthompson5910 (jim_thompson5910):

oh gotcha, yes the object speeds up on the interval (0,3.5) U (7.772, infinity). The object slows down on the interval (3.5, 7.772)

OpenStudy (babynini):

It's so gorgeous *-*

jimthompson5910 (jim_thompson5910):

why stop at 10? why not go off to infinity?

OpenStudy (babynini):

Right right.

OpenStudy (babynini):

Thank you so much

jimthompson5910 (jim_thompson5910):

you're welcome